### v2

# How simple can you go? An off-the-shelf transformer approach to molecular dynamics

Abstract
Most current neural networks for molecular dynamics (MD) include physical inductive biases, resulting in specialized and complex architectures. This is in contrast to most other machine learning domains, where specialist approaches are increasingly replaced by general-purpose architectures trained on vast datasets. In line with this trend, several recent studies have questioned the necessity of architectural features commonly found in MD models, such as built-in rotational equivariance or energy conservation. In this work, we contribute to the ongoing discussion by evaluating the performance of an MD model with as few specialized architectural features as possible. We present a recipe for MD using an Edge Transformer, an “off-the-shelf” transformer architecture that has been minimally modified for the MD domain, termed MD-ET. Our model implements neither built-in equivariance nor energy conservation. We use a simple supervised pre-training scheme on $\sim$30 million molecular structures from the QCML database. Using this “off-the-shelf” approach, we show state-of-the-art results on several benchmarks after fine-tuning for a small number of steps. Additionally, we examine the effects of being only approximately equivariant and energy conserving for MD simulations, proposing a novel method for distinguishing the errors resulting from non-equivariance from other sources of inaccuracies like numerical rounding errors. While our model exhibits runaway energy increases on larger structures, we show approximately energy-conserving NVE simulations for a range of small structures.
1 Introduction
Molecular dynamics (MD) simulations are essential for understanding molecular behavior [1, 2, 3], such as molecular relaxation [4], predicting structure [5], modeling interactions [6], replicating spectra [7], up to protein folding and design [8, 9, 10]. Machine Learning (ML) models can be used as heuristic approximations of the Schrödinger equation thereby significantly accelerating accurate MD simulations (e.g. Rupp et al. [11], Noé et al. [12], Unke et al. [13], Keith et al. [14], Unke et al. [15]).To ensure physical consistency of predictions, researchers have focused on incorporating inductive biases into MD models, which restrict models to physically plausible solutions [16, 17, 18, 19, 20, 13]. For instance, rotational equivariance guarantees that if a molecule is rotated, the forces are rotated as well (e.g. Schütt et al. [21], Bronstein et al. [22], Unke et al. [23]). Meanwhile, many other ML domains are increasingly using unconstrained, general-purpose architectures and large amounts of training data [24, 25, 26].
Several recent theoretical [27, 28] and empirical works [29, 30, 31, 32] discuss the necessity of commonly used physical constraints and inductive biases in MD models, most notably SO(3)-equivariance and energy conservation. For example, architectures predicting non-conservative forces include ForceNet [33], GemNet-dT [34], Orb [35] or Equiformer [36]. However, to the best of our knowledge, no architectures that fully embrace the general-purpose paradigm have been presented so far. We contribute by studying an extreme case: We present and evaluate the best MD model we can build using a minimal amount of MD-specific design features. Specifically, we aim to use an architecture that could be used in some other ML setting with minimal changes.
The “off-the-shelf” recipe we use is simple: We modify an Edge Transformer (ET) [37, 38] with MD-specific embedding layers and train on the new QCML database [39]. QCML has an unprecedented size and range across the periodic table, including out-of-equilibrium structures, different spin and charge states, and includes properties for a subset of $\sim$30 million entries calculated with density functional theory accuracy at the PBE0 level [40, 41] (including dispersion corrections [42, 43]). During training, we use two randomly rotated and mirrored copies of each sampled structure, so that our model can learn approximate O(3)-equivariance. Instead of predicting a combined loss [44] of energies and forces or calculating forces as the (negative) gradient of the predicted energy with respect to positions, we predict forces directly. This approach does not guarantee that the predicted forces are energy-conserving, but it improves the speed of the model, because forces need not be calculated by automatic differentiation. For downstream task evaluation, we fine-tune the model for a small number of steps only.
For MD simulations, we evaluate either the pre-trained model without further training (zero-shot) or fine-tune on a small number of samples of the target structure (few-shot). To increase the accuracy during MD simulations, we use a frame-averaging approach [45] over the SO(3) group to achieve an approximate “post-hoc” equivariance and further correct the predicted forces to guarantee no spurious rigid rotational or translational motion is introduced over time.
Our contributions are as follows:
  1. 1.
    We present an MD-adapted Edge Transformer (MD-ET) and show that our implementation can achieve competitive performance on several common benchmarks. We also demonstrate sample-efficiently finetuning.
  2. 2.
    We examine the effects of forgoing almost all commonly used inductive biases. Specifically, we evaluate the effects of being only approximately equivariant and energy-conserving in detail. We present a novel evaluation framework that is able to quantify deviations from true equivariance and distinguish them from effects stemming from model inaccuracy and numerical noise.
  3. 3.
    We present a comparison of MD-ET with an identically pretrained equivariant and energy-conserving model, highlighting the impact of removing physical constraints on MD simulations.
2 Background
While incorporating physics-inspired inductive biases into MLFF models yields powerful and data-efficient architectures, they also impose constraints. For instance, directly predicting (non-conservative) rotationally equivariant forces restricts a model to only performing equivariant operations [46]. A more subtle downside of specialized architectures is the increased difficulty to transfer advances from other ML research and profit from improved hardware and software tools. Engineering complexity was recently cited as the reason the AlphaFold project removed its equivariant inductive biases [10, 8]. Since protein conformer prediction is thematically close to MD, this prompted the discussion of which inductive biases are essential to the MD domain: Questioning the need for equivariance, Langer et al. [47] find that data augmentation during training can achieve a high degree of approximate equivariance in molecular dynamics. Neumann et al. [35] present Orb, a completely unconstrained message-passing architecture with outstanding benchmark performance. Elhag et al. [31] show that an additional loss term to induce equivariance combined with data augmentation results in competitive performance and approximate equivariance. Several theoretical works also suggest that unconstrained models can produce equivariant outputs under certain circumstances Gerken and Kessel [27], Nordenfors and Flinth [28], Puny et al. [45]. In contrast, the performance effects of removing energy conservation are less clear. Although benchmarking results suggest non-conservative models can perform stable MD simulations [35]. Bigi et al. [29] report that all non-conservative models in their study suffer from runaway temperature increases and thus instability.
Several current debates in the MD field center around competing paradigms – one leveraging inductive biases for sample efficiency, and another prioritizing simplicity, scale and transfer using general-purpose models. Recently, several aspects of the scaling paradigm have been studied and discussed in isolation [47, 29, 32]. Our work advances this discourse by empirically evaluating a simple baseline approach that fully adopts the scaling paradigm for MD and exploring its limits.
3 Molecular Dynamics Edge Transformer
We use the transformer, an architecture which has proven to scale well in many domains, as a starting point in our design process. We decide to use the Edge Transformer variant introduced by Bergen et al. [37] for algorithmic reasoning and adapted by Müller et al. [38] for graph learning, employing a higher-order attention mechanism instead of self-attention due to its expressivity (see below). We modify only the ET’s embeddings for MD. We therefore name our variant MD-ET.
Expressivity
Expressivity describes the ability of an ML model to tell the difference between two non-equivalent inputs, i.e., to distinguish two graphs that are not isomorphic. Morris et al. [48] show that the expressivity of a graph learning model can be related to performing a variant of the Weisfeiler-Leman (WL) graph isomorphism test. Hordan et al. [49] extend the WL-test to equivariant point clouds and show that a cutoff-free model with 3-WL expressivity is universal on 3D point clouds, i.e., such a model can theoretically learn to distinguish any two 3D point clouds. Müller et al. [38] show the ET has an expressive power equivalent to 3-WL. The ET can thus universally distinguish 3D point clouds, making it theoretically well-suited as an MLFF.
Tokenization
The ET derives its name from its tokenization scheme. We represent a molecular system as the set of all edges of a fully connected graph, that is, all pairs of atoms (including self-loops). We then encode the molecular system as a three-dimensional tensor $\boldsymbol{X}\in\mathbb{R}^{N\times N\times D}$, where $N$ denotes the number of atoms in the system and $D$ is the embedding dimension. An embedding $S_{ij}$ thus represents the edge between atoms $i$ and $j$, while $S_{ii}$ represents the atom $i$.
Triangular Attention Mechanism
To achieve the expressive power of 3-WL, the ET performs updates on the three-dimensional tensor representation described in the previous paragraph using a triangular attention mechanism, performing interactions between triples of atoms,
$\operatorname{TRIA}(\boldsymbol{x}_{ij})=\sum_{l=1}^{N}\alpha_{ilj}\boldsymbol{V}_{ilj}\ .$ (1)
Triangular attention is defined as a tensor product between an attention tensor $\boldsymbol{\alpha}\in\mathbb{R}^{N\times N\times N}$ and a value tensor $\boldsymbol{V}\in\mathbb{R}^{N\times N\times D}$. An element of the attention tensor is calculated as
$\alpha_{ilj}=\operatorname{softmax}_{l\in[N]}\left(\frac{1}{\sqrt{d}}\boldsymbol{x}_{il}\boldsymbol{W}^{Q}\left(\boldsymbol{x}_{lj}\boldsymbol{W}^{K}\right)^{\top}\right)\in\mathbb{R}\ ,$ (2)
i.e., the attention score between the representations of edges (i, l) and (l, j) and where the notation $[N]$ represents the set $\{1,2,\ldots,N\}$. The value vector
$\boldsymbol{v}_{ilj}=\boldsymbol{x}_{il}\boldsymbol{W}^{V_{1}}\odot\boldsymbol{x}_{lj}\boldsymbol{W}^{V_{2}}\in\mathbb{R}^{D}$ (3)
is a combination of the value vectors of $\boldsymbol{x}_{il}$ and $\boldsymbol{x}_{lj}$ through elementwise multiplication. $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$,$\boldsymbol{W}^{V_{1}}$, $\boldsymbol{W}^{V_{2}}$ $\in\mathbb{R}^{d\times d}$ are learned weight matrices [38]. The triangular attention mechanism extends to multi-head attention analogous to regular self-attention. Similarly, a full ET layer is defined analogously to a transformer layer [50]; specifically,
$\boldsymbol{x}^{(t)}_{ij}=\operatorname{FFN}\left(\operatorname{TRIA}\left(\operatorname{LN}\left(\boldsymbol{x}^{(t-1)}_{ij}\right)\right)+\boldsymbol{x}^{(t-1)}_{ij}\right),$ (4)
where $\operatorname{FFN}$ is a feed-forward neural network and $\operatorname{LN}$ denotes layer norm [51].
The triangular attention is straightforward to implement and JIT-compile (see Appendix A.2). The large 3D tensor products are embarrassingly parallelizable and thus execute very efficiently on accelerators. This makes the ET comparatively fast despite a runtime and memory complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system (see Appendix A.3).
Molecular Embeddings
To generate initial edge representations, we combine several embeddings: i) spin and charge, ii) atomic numbers , iii) pairwise distances, and iv) pairwise displacement vectors. To maintain MD-ET’s simplicity we implement commonly used embedding layer where possible. For details, see Appendix A.2.1.
4 Experiments
To contribute to the current discussion about inductive biases in MD, we need to thoroughly evaluate MD-ET. Specifically, we want to address the following questions:
  • Q1: Can an “off-the-shelf” model without inductive biases compete on common benchmarks measuring accuracy, inference speed, and MD stability?
  • Q2: How equivariant is MD-ET? Is approximate equivariance likely to impact MD simulation stability?
  • Q3: Is MD-ET approximately energy-conserving? How affected is the stability of MD simulations?
  • Q4: How faithful are MD simulations? Is there a correlation between energy conservation and observables?
4.1 Pretraining
We use a pretrained model for all downstream evaluations. However, instead of the commonly used denoising task [52, 35] we perform supervised pretraining on the new QCML dataset [39]. We create an approximate 90%/5%/5% split from QCML. As QCML contains multiple conformations for each structure sampled along their normal modes, we ensure that all conformations of a structure get assigned to the same split.
To learn approximate equivariance, we use data augmentation: During pretraining we duplicate each batch once and apply random rotations and reflections to both copies to form an augmented batch. We train MD-ET for 880k steps on QCML using a batch size of 1024 (i.e., 512 before data augmentation), which takes approximately 16 A100 GPU-days (see Appendix A.5).
4.2 Postprocessing
Our model does not require any modifications for downstream fine-tuning tasks. For MD-simulations we apply two techniques to enhance model predictions:
Net Force and Torque Removal
We adopt the net force and torque removal algorithm described in [35]. We find that removing net forces and torque from model predictions improves the stability of MD simulations, while only increasing the time of a model evaluation by 10-20 percent (see Appendix A.3). We use net force and torque removal in all simulation experiments.
Frame-averaging
For some simulations, we observe that frame-averaging [45] helps to improve MD stability. For frame-averaging we use $M$ rotation matrices to pass $M$ different orientations of the input positions to the model. We then apply the inverse rotations to the predicted forces and average them. We discuss the effects of frame-averaging in section 4.4. We use frame-averaging only where explicitly stated. For implementation details, see section 4.4 and Appendix A.5.
4.3 Q1: Benchmark Results
We evaluate the performance of our proposed MD-ET model across several standard molecular dynamics benchmarks. We compare MD-ET with established MLFFs, focusing on their accuracy, stability and efficiency. Our experiments cover a range of datasets, including QCML [39], MD17 [16], and Ko2020 [53], each presenting distinct challenges.
Performance on the QCML Dataset
Table 1: Force prediction test MAE on the QCML dataset in kcal/mol Å-1 for different models.
Model Test MAE
MD-ET 0.69
SpookyNet 0.74
PaiNN 0.98
Does not separate conformers between training and testing.
In Table 1, we present the mean absolute error (MAE) of force predictions for MD-ET, SpookyNet [13], and PaiNN [21].
MD-ET achieves the lowest MAE among the compared models, outperforming both SpookyNet and PaiNN. We observe that SpookyNet’s results do not separate conformers between training and testing, which may overestimate generalization performance. This improvement over equivariant models like PaiNN and SpookyNet suggests that approximate equivariance, as utilized in MD-ET, does not hinder performance on such benchmarks. However, a low test error does not necessarily imply a low MD simulation error or a stable trajectory, as we will demonstrate below.
MD17-10k: molecule specific results
Table 2: Performance comparison on the MD17 benchmark on stability (ps) with a maximum at 300 ps, force MAE (meV Å-1) and FPS. Higher is better is denoted ()($\uparrow$)( ↑ ) and vice versa. Best average results in bold. Results marked with a indicate that the molecule was included in the pretrain train split.
Metric Molecule DeepPot-SE [54] SchNet [17] DimeNet [55] PaiNN [21] SphereNet [56] ForceNet [33] GemNet-dT [34] NequIP [57] MD-ET (pretrained) (ours) MD-ET (finetuned) (ours)
Stability ($\uparrow$) Aspirin 9 26 54 159 141 182 192 300 154 300
Ethanol 300 247 26 86 33 300 300 300 300 300
Naphthalene 246 18 85 300 6 300 25 300 256 300
Salicylic Acid 300 300 73 281 36 1 94 300 213 300
Average 213.8 147.8 60.0 206.5 54 195.8 152.8 300 230.8 300
MAE ($\downarrow$) Aspirin 21.0 35.6 10.0 9.2 3.4 22.1 5.1 2.3 4.2 4.2
Ethanol 8.9 16.8 4.2 5.0 1.7 14.9 1.7 1.3 2.5 $1.0^{*}$
Naphthalene 13.4 22.5 5.7 3.8 1.5 9.9 1.9 1.1 4.0 2.3
Salicylic Acid 14.9 26.3 9.6 6.5 2.6 12.8 4.0 1.6 3.3 2.8
Average 14.6 25.3 7.4 6.1 2.3 15.0 3.2 1.6 3.5 2.6
FPS ($\uparrow$) Aspirin 88.0 108.9 20.6 85.8 17.5 137.3 56.8 8.4 159.8 159.8
Ethanol 101.0 112.6 21.4 87.3 30.5 141.1 54.3 8.9 148.1 148.1
Naphthalene 109.3 110.9 19.1 92.8 18.3 140.2 53.5 8.2 150.2 150.2
Salicylic Acid 94.6 111.7 19.4 90.5 21.4 143.2 52.4 8.4 147.3 147.3
Average 98.2 111.0 20.1 89.1 21.9 140.5 54.3 8.5 151.4 151.4
The MD17 dataset [16] consists of molecular dynamics trajectories for small organic molecules and has been proposed as a benchmark by Fu et al. [58], challenging models to not only predict a low force error but also measures stability, i.e., how long a simulation experiences no catastrophic failure. The third criterion is speed, i.e., how many times per second the model can be evaluated. Table 2 compares MD-ET with several state-of-the-art models, including NequIP [57], GemNet-dT [34]. In addition to training ET-MD directly on MD17, we also report fine-tuning results after 2000 fine-tuning steps (see Appendix A.6 for details). Since there is some overlap with the pretraining set, fine-tuning results for ethanol (marked with a *), have only limited significance.
Table 3: MAE of atomic forces (meV Å-1) for MD17 molecules recalculated with CCSD(T). Best results in bold. Molecules with a are included in the pretrain dataset of MD-ET.
Molecule sGDML [20] GemNet-Q [34] GemNet-T [34] NewtonNet [59] NequIP [57] MD-ET (1000 samples) (ours, finetuned) MD-ET (100 samples) (ours, finetuned)
Ethanol 15.0164 3.0814 3.0814 10.2424 2.9946 1.4683 $\pm$ 0.0017 $2.5756\pm 0.0045$
Aspirin 33.0274 10.4160 10.3292 15.4504 8.2894 4.9525 $\pm$ 0.0069 $9.0772\pm 1.0465$
Benzene 1.8228 0.6944 0.6944 0.4774 0.2604 $2.0507\pm 0.0645$ $2.7354\pm 0.0123$
Malonaldehyde 16.2316 4.3400 5.9024 12.3690 4.5136 2.5632 $\pm$ 0.0158 $5.1715\pm 0.0688$
Toluene 8.8970 2.5172 2.6908 3.4720 1.6926 $2.6117\pm 0.0072$ $4.3461\pm 0.0062$
MD-ET achieves force MAEs similar to those of NequIP and GemNet-dT (which is also not energy-conserving). The only other model that violates energy conservation is ForceNet at a much higher MAE. While NequIP attains slightly lower MAEs, MD-ET excels in computational efficiency, processing over 150 frames per second (FPS) beating all other models. MD-ET also reaches the maximum simulation stability score across all molecules. While a 300 ps NVT simulation is not sufficient to assess the long-term stability of MD trajectories, it can still be used to compare the stability of different models.
MD17@CCSD(T)
For molecular dynamics (MD) simulations where precise electron correlation is necessary to accurately model potential energy surfaces and critical interactions, such as transition states, Density Functional Theory (DFT) reference data may not always be sufficiently accurate. The CCSD(T) level of theory is often regarded as the “gold standard” of quantum chemistry and is thus a suitable choice for these cases. Nonetheless, the substantial computational cost of CCSD(T) renders it impractical for large-scale or long-duration simulations. Transfer learning from lower-level theories, such as DFT, presents an effective strategy to approximate CCSD(T) using the expensive CCSD(T) samples efficiently. We use the MD17@CCSD(T) dataset [20], which recalculates trajectories for molecules in MD17 using CCSD(T) to test this approach using MD-ET. Although MD-ET was pretrained on the PBE0 level, Table 3 demonstrates that MD-ET can effectively capture higher-level interactions through fine-tuning with a limited number of CCSD(T) samples. When fine-tuned on 1,000 CCSD(T)-level samples, MD-ET achieves lower mean absolute error (MAE) than all competing models for three of the five structures. For toluene, its MAE is comparable to GemNet-T, while for benzene, it exhibits the highest error among the benchmarked models. Furthermore, when we reduce the fine-tuning dataset by an order of magnitude MD-ET still reaches accuracy comparable to models trained on the full dataset.
Ko2020: molecule specific results
Ko et al. [53] introduce a molecular dataset consisting of molecular (e.g. $\text{C}_{10}\text{H}_{2}$ and $\text{C}_{10}\text{H}^{+}$) and metallic systems (e.g., $\text{Ag}_{3}^{+/-}$), which are typically challenging for conventional MLFFs. Table 4 compares the force root mean squared errors (RMSE) of MD-ET with 4G-BPNN [53] and SpookyNet [13] on selected systems from the Ko2020 dataset.
Table 4: RMSEs of atomic forces (meV Å-1) for different molecular systems, comparing 4G-BPNN, SpookyNet, and our work. The values for 4G-BPNN are taken from [53]. Best results in bold.
Molecule 4G-BPNN [53] SpookyNet [13] MD-ET (ours, finetuned)
C10H2/C10H+${}_{3}^{+}$ 78.00 5.802 5.122 $\pm$ 0.197
Na8/Cl+${}_{8}^{+}$ 32.78 1.052 2.26 $\pm$ 0.014
Ag+/${}_{3}^{+/-}$ 31.69 26.64 11.78 $\pm$ 0.409
Au2–MgO 66.0 5.337 -
For the C10H2/C10H+${}_{3}^{+}$ and Ag+/${}_{3}^{+/-}$ systems, MD-ET outperforms both 4G-BPNN and SpookyNet. This demonstrates MD-ET’s capability to model smaller ionic and metallic systems. However, on the Na8/Cl+${}_{8}^{+}$ MD-ET’s performance is slightly worse and fails to converge on Au2–MgO systems.
4.4 Q2: Equivariance Evaluation
Figure 1: Equivariance Evaluation. Top Left: Average equivariance error $E_{\text{eq}}$ for 2048 randomly sampled QCML train and test structures using different frame-averaging sample sizes $n$. Top Right: Equivariance error (without frame averaging) over the $\mathrm{SO(3)}$ group across a range of alkanes and cumulenes. Bottom Left: Approximate numerical noise level over the $\mathrm{SO(3)}$ group for alkanes and cumulenes. Bottom Right: Relative equivariance error for alkanes and cumulenes.
Our model, MD-ET, is only approximately equivariant. Although MD-ET’s performs well on benchmarks, we examine the potential influence of approximate equivariance on MD simulations more closely. To quantify how (non-)equivariant our model is, we define the equivariance error as the expected Euclidean distance between the model’s prediction on a randomly rotated input and the expected prediction averaged over all rotations in the symmetry group,
$E_{\text{eq}}({D},\hat{f}_{\theta})=\underset{\begin{subarray}{c}x\sim{D}\\ \mathcal{S}\sim\text{SO}(3)\end{subarray}}{\mathbb{E}}\left[\left\|\underset{\mathcal{R}\sim\text{SO}(3)}{\mathbb{E}}\left[\mathcal{R}^{\top}\hat{f}_{\theta}(\mathcal{R}{x})\right]-\mathcal{S}^{\top}\hat{f}_{\theta}(\mathcal{S}{x})\right\|_{2}\right],$ (5)
where $x\sim{D}$ are the 3D positions sampled from their distribution $D$, $\mathcal{R},\mathcal{S}\sim\text{SO}(3)$ are rotation matrices, and $\hat{f}_{\theta}$ is a trained predictor. For a perfectly equivariant model, $E_{\text{eq}}=0$ (proof in Appendix A.4).
In practice, models use floating point arithmetic with finite precision and numerical inaccuracies accumulate during evaluation. For this reason, equivariant models do not reach $E_{\text{eq}}=0$. To better assess MD-ET’s deviation from an equivariant architecture, we evaluate $E_{\text{eq}}$ using different amounts of frame-averaging (see section 4.2). Concretely, this means averaging the model predictions over $n$ randomly sampled rotations $\mathcal{R}_{i}\in\mathrm{SO(3)}$,
$\hat{f}_{\theta}^{\prime}(x)=\int_{\mathrm{SO(3)}}\mathcal{R}^{\top}\hat{f}_{\theta}(\mathcal{R}x)\,\mathrm{d}\mu(\mathcal{R})\approx\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}x).$ (6)
The limit, $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},f^{\prime})$, converges to zero for any $f_{\theta}^{\prime}$ (proof in Appendix A.4). When empirically evaluating $E_{\text{eq}}(\mathcal{D},\hat{f}_{\theta}^{\prime})$ for high values of $n$, we can thus attribute remaining errors to numerical inaccuracies during evaluation. We can thus grasp the numerical limit of $\mathrm{SO}(3)$ equivariance and how close our model has approached it. In practice we approximate $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},f^{\prime})$ by evaluating $E_{\text{eq}}$ for multiple frame-averaging sizes $n$ and fitting an exponential function to extrapolate to the asymptotic value for $n\to\infty$.
We evaluate MD-ET’s average $E_{\text{eq}}$ across 2048 randomly drawn QCML train and test samples (see Figure 1, top left). The results show that without frame-averaging MD-ET reaches an $E_{\text{eq}}$ of around $0.06$ kcal/mol/Å (top left), which is one order of magnitude below our test set loss and two to three orders of magnitude below typical force magnitudes. Frame-averaging exponentially decreases the equivariance error towards a value below $0.02$ kcal/mol/Å (bottom right). $E_{\text{eq}}$ is higher on the test set, which suggests that learned equivariance might not be retained further from the training distribution. To investigate this possibility, we calculate $E_{\text{eq}}$ for a number of alkanes (purple) and cumulenes (blue) of increasing size: Structures with more than 8 carbons are out-of-distribution (see Figure 1, top right; for visualization, see Appendix B). For each structure, we approximate the numerical noise level (see Figure 1, bottom left) using MD-ET and an equivariant SpookyNet. We also examine the ratio between the noise level and $E_{\text{eq}}$ without frame-averaging (see Figure 1, bottom right) to evaluate how closely MD-ET approaches the practical limit for SO(3) equivariance across structures.
Results demonstrate that MD-ET’s approximation of SO(3) equivariance measured relative to the noise level varies little between structures, even in the extrapolative regime. On the other hand, numerical noise varies significantly between structures and tends to increase with structure size. This observation extends to the equivariant SpookyNet architecture.
For in-distribution data, we find that MD-ET learns to approximate equivariance well, approaching the limit imposed by the numerical accuracy to an order of magnitude. Given that we use only one additional data augmentation per sample during training, learning approximate equivariance appears cheap and effective. The achieved equivariance error is unlikely to impact force prediction accuracy since it is two orders of magnitude lower than typical molecular forces for most structures. Four frame-averaging samples can further reduce the equivariance error by more than 50 %. In the extrapolative regime, we observe an increase in numerical noise and larger differences in noise levels between structures. While problems with out-of-distribution generalization are widely studied, deteriorating numerical stability in the extrapolative regime is a distinct phenomenon and has a significant effect on an MLFF’s prediction quality. In contrast, learned equivariance appears to generalize well when noise levels are taken into account. We thus conclude that learning approximate equivariance is unlikely to impede a model’s accuracy, while numerical instability–which is a concern for all ML models–could.
4.5 Q3: Evaluation of Energy Conservation
Figure 2: Total energy variation over time for NVE simulations of linear alkanes CnH2n+2 ($n=1\dots 8$). For reference, we compare to MD simulations with an energy-conserving model (SpookyNet [13]) trained on the same data as MD-ET. A linear fit is plotted above the raw data to help visualize energy drift.
Since MD-ET directly predicts forces without enforcing energy conservation, we aim to assess whether the predicted forces are approximately conservative in addition to being approximately equivariant. Although MD-ET achieves stable MD simulations on the MD17 benchmark (see Table 2), this alone is not sufficient to demonstrate approximate energy conservation. This is because the dynamics sample the canonical (NVT) ensemble, i.e., the aim is to keep the temperature of the simulated system constant. To achieve this, a thermostat constantly introduces or removes energy, which prevents testing energy conservation. We therefore also perform MD simulations sampling the microcanonical (NVE) ensemble, where total energy is a conserved quantity. This enables quantification of how strongly a simulation violates energy conservation by monitoring the variation of total energy over time. Note that even when using conservative FFs, the total energy typically fluctuates slightly due to numerical noise and discretization errors when integrating the equations of motion. This is acceptable as long as these fluctuations are small and stay centered around zero, i.e., do not introduce energy drift. We find that NVE simulations with MD-ET are approximately energy-conserving for small systems (fewer than around 15 atoms), but for larger systems, the total energy tends to increase linearly with time (see Figure 2). The additional energy causes increasingly large structural fluctuations until the simulation invariably becomes unstable after a sufficiently large number of time steps. For many MD experiments MD-ET’s utility is thus questionable. Nevertheless, our results are surprising: To our knowledge no other approximately energy-conserving MD model has been shown to be approximately energy-conserving and stable in short NVE simulations [29].
4.6 Q4: MD Simulation results
Figure 3: Left column: Dihedral angle of cumulenes of length 6 to 9 over an NVT simulation of 30 ps. Right column: Instantaneous temperature over time. The inset in the first two rows shows the flat structure of the first cumulene and the perpendicular structure in the second one. The dihedral angle $\omega$ is denoted in red, encompassing the four hydrogen atoms.
[…]
5 Conclusion
There is an ongoing debate in the MD community whether including inductive biases such as energy conservation or rotational equivariance in ML architectures are necessary for successful MD applications. ML methods have always needed to find a balance between ease of optimization and the need to include helpful inductive biases. While traditional models for MD lean toward strong inductive biases, here, we present MD-ET, an Edge Transformer-based approach to molecular dynamics simulations that relaxes strict equivariance and energy conservation constraints during training. By predicting forces directly and utilizing data augmentation, our model achieves approximate equivariance and is approximately energy-conserving for MD simulations of small systems. MD-ET is fast, taking full advantage of the accelerators it runs on. It closely resembles a regular transformer, uses no custom layers (not counting embedding layers) and is easy to implement and use. By being “off-the-shelf”, MD-ET can profit from a vast body of research on transformers and transformer variants. It is uncomplicated to train and can quickly be applied to new datasets.
For a minimally MD-adapted architecture, MD-ET shows surprisingly good benchmark results, beating specialized architectures on several tasks measuring accuracy, speed and stability. It further displays good few-shot transfer and zero-shot capabilities, including stable zero-shot short MD simulations in the NVT ensemble. Longer simulations, especially in the NVE ensemble, exhibit instabilities. Such problems are also known for equivariant and energy-conserving models, where they are often associated with so-called “holes in the potential energy surface” [60].
To better understand potential problems with unconstrained architectures we experimented with MD-ET, studying potentially problematic properties. To evaluate equivariance we introduced a simple metric which is able to usefully quantify the expected error incurred by approximate equivariance and isolate it from other sources of error such as numeric instability. Results suggest that numeric instability for MD-models in general significantly varies from structure to structure and grows for larger structures and in the extrapolative regime. With regards to learning approximate energy conservation, we find that, at least with our current approach, this seems to work reasonably well for small systems of less than 15 atoms. Here too, larger systems prove more challenging.
MD-ET’s performance suggests that models for MD simulations may not necessarily need a special-purpose architecture to adhere to physical constraints, and instead, these constraints can also be learned from data. This is particularly successful for rotational equivariance, which MD-ET can learn almost up to numerical precision. Whether energy conservation can similarly be learned and whether non-conservative MLFFs can be used for reliable MD simulations of large molecular systems is questionable, at least without further advances, such as additional loss terms that encourage learning of energy-conserving forces. The disagreement of metrics commonly used by the ML community and MD simulation results is not an issue limited to MD-ET but instead raises the question how well ML benchmarks are aligned with the needs of practitioners.
Our experiments with MD-ET show an uncommon and unexpectedly successful new approach to MD (for summary see Appendix A.1). While many problems remain and unconstrained architectures such as MD-ET are still unsuitable for many MD simulations, we have promising early results. We have also identified several avenues of further investigation such as length generalization ability and the impact of numerical noise on generalization.
Software and Data
Project code is available on github: https://github.com/mx-e/simple-md.
Acknowledgements
We thank Luis Müller for his encouragement and being a great colleague. We also thank Elron Pens, Stefan Chmiela and Sidney Bender for important insights and fruitful discussions. Stefan Gugler was supported by the Postdoc.Mobility fellowship by the Swiss National Science Foundation (project no. 225476)
Appendix A Appendix
A.1 Side-by-side comparison
Table 5: Comparison of different force field approaches: MD-ET, MLFF (typical learned force fields), and Traditional FF (classical potentials).
Feature MD-ET MLFF Traditional FF
Equivariance (close to numerical precision)
Energy Conservation
Short NVE Stability $\sim$
Long NVE Stability
Short NVT Stability
Long NVT Stability $\sim$
Differentiable
Fast (per step)
Accurate (quantum-level)
Low Inductive Bias $\sim$
A.2 Edge Transformer Implementation
A.2.1 Molecular embedding layer details
To create initial edge representations, we combine several embeddings: i) spin & charge, ii) atomic numbers , iii) pairwise distances, and iv) pairwise directions.
Spin and charge are embedded through learned embedding layers. The total spin, $s_{i}\in\mathbb{N}_{0}$, denotes the number of unpaired electrons, and $q_{i}\in\mathbb{Z}$ denotes the net charge. They are jointly embedded as
$\mathbf{z}_{i}=\mathbf{W}^{\text{spin}}{h}(s_{i})+\mathbf{W}^{\text{charge}}{h}(q_{i}),$ (7)
where $\mathbf{W}^{\text{spin}}\in\mathbb{R}^{D\times d_{\text{spin}}}$ and $\mathbf{W}^{\text{charge}}\in\mathbb{R}^{D\times d_{\text{charge}}}$ are learned weight matrices for spin and charge, respectively, ${h}$ is a one-hot encoding and $d_{\text{spin}}$ and $d_{\text{charge}}$ are the sizes of the spin and charge vocabularies, respectively.
The edge embedding for atoms $i$ and $j$ is defined as
$A_{ij}=\psi_{\text{edge}}\Big{(}$ $\left[\mathbf{W}^{\text{atom}}{h}(Z_{i})+\mathbf{W}^{\text{cfg}}\tau(Z_{i})+\mathbf{z}_{i},\right.$
$\left.\mathbf{W}^{\text{atom}}{h}(Z_{j})+\mathbf{W}^{\text{cfg}}\tau(Z_{j})+\mathbf{z}_{j}\right]\Big{)}\ ,$ (8)
where $Z_{i},Z_{j}\in\mathbb{N}$ are the atomic numbers and $\mathbf{W}^{\text{atom}}\in\mathbb{R}^{D\times d_{\text{atom}}}$ and $\mathbf{W}^{\text{cfg}}\in\mathbb{R}^{D\times d_{\text{cfg}}}$ are the learned weight matrices for atomic numbers and electron configuration, $\tau$ (see Appendix A.2.1), respectively. Lastly, $\psi_{\text{edge}}:\mathbb{R}^{2D}\rightarrow\mathbb{R}^{D}$ is an MLP that projects concatenated embeddings into the embedding space.
To embed the pairwise distance between atoms we use a set of $K=128$ radial basis functions (RBF),
$\varphi_{k}(x)=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\left(-\frac{(x-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right)$ (9)
with learnable parameters $\mu_{k}$ and $\sigma_{k}$, where the former is initialized on $\mathcal{U}(0,7)$ and the latter on $\mathcal{U}(0,3)$, to obtain
$R_{ij}=\psi_{\text{dist}}\left(\sum_{k=1}^{K}\varphi_{k}(md_{ij}+b)\right),$ (10)
where $\psi_{\text{dist}}$ is an MLP and $m$, $b$, are also learnable parameters initialized as 1 and 0, respectively and $d_{ij}$ being the interatomic distance.
Directional embeddings are created by embedding the azimuthal and polar angles $\phi$ and $\theta$ of the pairwise displacement vectors using $K^{\prime}=128$ Fourier kernels and a linear layer. The azimuthal frequencies and polar frequencies are defined as
$\omega_{\phi}^{k^{\prime}}=\pi\left(\frac{1}{2\pi}\right)^{\frac{k^{\prime}}{K^{\prime}/2-1}}\quad\text{and}\quad\omega_{\theta}^{k^{\prime}}=\pi\left(\frac{1}{\pi}\right)^{\frac{k^{\prime}}{K^{\prime}/2-1}},$ (11)
which covers the angular space for a sufficient $K^{\prime}$. With the Fourier feature matrix
$F(\phi,\theta)=\begin{bmatrix}\sin(\phi\boldsymbol{\omega}_{\phi})&\cos(\phi\boldsymbol{\omega}_{\phi})&\sin(\theta\boldsymbol{\omega}_{\theta})&\cos(\theta\boldsymbol{\omega}_{\theta})\end{bmatrix}$ (12)
the directional embedding is
$\Gamma_{ij}=\psi_{\text{dir}}\left(F(\phi,\theta)\right),$ (13)
where $\psi_{\text{dir}}$ is another learnable linear layer.
Finally, all embeddings are added so the final edge embedding is
$E_{ij}=A_{ij}+R_{ij}+\Gamma_{ij}\,.$ (14)
Prediction Head & Training Objective
Forces are predicted directly, i.e.,
${\boldsymbol{\hat{f}}}_{i}=\psi_{3}\left(\sum^{N}_{l=1}\psi_{1}(\boldsymbol{x}_{il})+\psi_{2}(\boldsymbol{x}_{lj})\right)\in\mathbb{R}^{3},$ (15)
where $\psi_{\{1:3\}}$ are MLPs. The loss functional for training is
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\|\boldsymbol{\hat{f}}_{i}-\boldsymbol{f}_{i}^{*}\|_{2}\ ,$ (16)
where $\|\cdot,\cdot\|_{2}$ is the $\ell^{2}$ norm between prediction and the ground truth, ${\boldsymbol{f}}_{i}^{*}$, and $N$ is the number of atoms in the structure.
Electronic embedding
The function $\tau:\mathbb{N}\rightarrow\mathbb{R}^{d_{\text{cfg}}}$ maps an atomic number $Z_{i}$ to a fixed electron configuration descriptor that encodes i) subshell occupancy, the number of electrons in each $s$, $p$, $d$, and $f$ subshell up to the valence shell, ii) valence electrons in the outermost $s$ (vs), $p$ (vp), $d$ (vd), and $f$ (vf) subshells.
This descriptor captures an atom’s electronic structure, which governs its chemical behavior. For example, chlorine ($Z=17$) and bromine ($Z=35$) both have $7$ valence electrons ($\text{vs}=2$, $\text{vp}=5$), explaining their tendency to form -1 ions. The configuration follows the Aufbau principle, with exceptions in transition metals and lanthanides/actinides accounted for empirically.
$\tau(Z_{i})$ retrieves a precomputed vector from a lookup table (excerpt below). Each row corresponds to an element’s ground-state electron configuration and valence electrons. The full table spans $Z=1$ (H) to $Z=100$ (Fm).
Table 6: Electron configuration descriptor examples. Full table available in source code.
Z Element 1s 2s 2p 3s 3p vs/vp/vd/vf
1 H 1 0 0 0 0 1/0/0/0
8 O 2 2 4 0 0 2/4/0/0
17 Cl 2 2 6 2 5 2/5/0/0
35 Br 2 2 6 2 6 2/5/10/0
79 Au 2 2 6 2 6 1/0/10/14
The descriptor vector is linearly transformed by $\mathbf{W}^{\text{cfg}}\in\mathbb{R}^{D\times d_{\text{cfg}}}$ to produce a $D$-dimensional embedding. This allows the model to learn similarities between elements with analogous configurations (e.g., halogens, noble gases) directly from their electronic structure.
ET Layer implementation details
We implement MD-ET using PyTorch 2.5.1. Pseudo-code of the triangular and a comparison with regular (single-head) attention can be found in Algorithm 1 (below). To improve performance, we compile the triangular attention operation at runtime using torch compile. This necessitates rewriting triangular attention without einsums.
A.3 Inference Speed Evaluation
See Table 7 below. For runtime benchmarks, we time the forward pass of the model, including all postprocessing steps on a cloud-based V100 GPU (Throughput, Latency). Inference Time measures the pure inference time without postprocessing. To give an approximation of the performance impact of frame-averaging we report the model throughput for multiple batch sizes and structures of different sizes. To make the results more comparable to different hardware we also provide total FLOPs per forward pass.
$\begin{array}[]{ccccccc}\hline\cr\hline\cr\text{Molecule}&\text{Batch Size}&\text{Throughput}&\text{Avg Latency (ms)}&\text{Inference Time (ms)}&\text{FLOPs}\\ \hline\cr\text{Aspirin}&1&159.800&6.258&4.927&5.42\times 10^{9}\\ &2&157.909&6.333&5.063&1.08\times 10^{10}\\ &4&147.431&6.783&5.505&2.17\times 10^{10}\\ &8&102.680&9.739&8.458&4.33\times 10^{10}\\ &16&63.636&15.714&14.414&8.67\times 10^{10}\\ &32&35.745&27.976&26.652&1.73\times 10^{11}\\ &64&18.685&53.520&52.103&3.47\times 10^{11}\\ \hline\cr\text{Ethanol}&1&148.151&6.750&5.429&9.96\times 10^{8}\\ &2&146.148&6.842&5.586&1.99\times 10^{9}\\ &4&151.253&6.611&5.343&3.98\times 10^{9}\\ &8&153.357&6.521&5.247&7.97\times 10^{9}\\ &16&151.867&6.585&5.313&1.59\times 10^{10}\\ &32&115.476&8.660&7.377&3.19\times 10^{10}\\ &64&70.741&14.136&12.784&6.37\times 10^{10}\\ \hline\cr\text{Naphthalene}&1&150.254&6.655&5.320&3.98\times 10^{9}\\ &2&151.231&6.612&5.333&7.96\times 10^{9}\\ &4&151.409&6.605&5.332&1.59\times 10^{10}\\ &8&114.743&8.715&7.458&3.18\times 10^{10}\\ &16&70.293&14.226&12.937&6.37\times 10^{10}\\ &32&41.122&24.318&22.987&1.27\times 10^{11}\\ &64&22.033&45.387&43.976&2.55\times 10^{11}\\ \hline\cr\text{Salicylic Acid}&1&147.294&6.789&5.481&3.14\times 10^{9}\\ &2&150.733&6.634&5.371&6.29\times 10^{9}\\ &4&154.547&6.471&5.225&1.26\times 10^{10}\\ &8&130.698&7.651&6.407&2.52\times 10^{10}\\ &16&85.649&11.676&10.401&5.03\times 10^{10}\\ &32&51.222&19.523&18.236&1.01\times 10^{11}\\ &64&27.834&35.927&34.564&2.01\times 10^{11}\\ \hline\cr\hline\cr\end{array}$
Table 7: Inference Speed Evaluation
A.4 Equivariance proof
See main text for the equivariance error, Eq. (5).
Proof: We first simplify the first term and substitute the equivariance property $f_{\theta}(\mathcal{R}\boldsymbol{x})=\mathcal{R}f_{\theta}(\boldsymbol{x})$:
$\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}f_{\theta}(\mathcal{R}\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}\mathcal{R}f_{\theta}(\boldsymbol{x})\right].$ (17)
Since $\mathcal{R}^{\top}\mathcal{R}=\boldsymbol{I}$ (identity matrix) for all $\mathcal{R}\in\text{SO}(3)$:
$\mathbb{E}_{\mathcal{R}}\left[\boldsymbol{I}f_{\theta}(\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[f_{\theta}(\boldsymbol{x})\right].$ (18)
The expectation $\mathbb{E}_{\mathcal{R}}[f_{\theta}(\boldsymbol{x})]$ is over $\mathcal{R}$, but $f_{\theta}(\boldsymbol{x})$ does not depend on $\mathcal{R}$. So we obtain
$\mathbb{E}_{\mathcal{R}}\left[f_{\theta}(\boldsymbol{x})\right]=f_{\theta}(\boldsymbol{x}).$ (19)
For the 2nd term we apply the equivariance property to $f_{\theta}(\mathcal{S}\boldsymbol{x})$
$f_{\theta}(\mathcal{S}\boldsymbol{x})=\mathcal{S}f_{\theta}(\boldsymbol{x}).$ (20)
and similarly obtain
$\mathcal{S}^{\top}f_{\theta}(\mathcal{S}\boldsymbol{x})=\mathcal{S}^{\top}\mathcal{S}f_{\theta}(\boldsymbol{x})=\boldsymbol{I}f_{\theta}(\boldsymbol{x})=f_{\theta}(\boldsymbol{x}).$ (21)
Therefore, the entire expectation over $x$ and $\mathcal{S}$ is
$\mathbb{E}_{x,\mathcal{S}}[||f_{\theta}(\boldsymbol{x})-f_{\theta}(\boldsymbol{x})||]=0.$ (22)
2nd Claim: $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},\hat{f}^{\prime})=0$ for all neural nets.
Proof: We insert $\hat{f_{\theta}}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}x)$ into $\mathcal{S}^{\top}f_{\theta}(\mathcal{S}\boldsymbol{x})$, obtaining:
$\mathcal{S}^{\top}\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}\mathcal{S}\boldsymbol{x})$ (23)
Rearranging, we obtain:
$\frac{1}{n}\sum_{i=1}^{n}(\mathcal{R}_{i}\mathcal{S})^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}\mathcal{S}\boldsymbol{x})$ (24)
We now substitute $\mathcal{R}_{i}\mathcal{S}=\mathcal{T}_{i}\in SO(3)$, obtaining
$\frac{1}{n}\sum_{i=1}^{n}(\mathcal{T}_{i})^{\top}\hat{f}_{\theta}(\mathcal{T}_{i}\boldsymbol{x})$ (25)
According to Kolmogorov’s law the sample mean converges to the expectation for large numbers of samples. For $\lim_{n\rightarrow\infty}$ We therefore obtain
$\mathbb{E}_{\mathcal{T}\sim SO(3)}[\mathcal{T}^{\top}\hat{f}_{\theta}(\mathcal{T}\boldsymbol{x})]$ (26)
Inserting into equation 5 it is trivial to see that $E_{\text{eq}}(\mathcal{D},\hat{f}^{\prime})=0$.
A.5 Training Protocol
MD-ET was trained on the QCML dataset. We employed a cosine warmup learning rate schedule with an initial warmup phase of 5,000 steps, starting from 1e-6 and peaking at 5e-4. We conducted training on 8 NVIDIA A100 GPUs with 80GB memory each. Important hyperparameters include:
Parameter Value
Learning Rate $5\times 10^{-4}$
Batch Size 1024
Total Steps 880,000
Weight Decay $1\times 10^{-7}$
Gradient Clip 1.0
Embedding Dimension 192
Number of Layers 12
Number of Heads 12
FFN Multiplier 4
3D Kernels 128
Warmup Steps 5000
EMA Decay 0.9997
Minimum Learning Rate $5\times 10^{-8}$
Attention Dropout 0
FFN Dropout 0
Table 8: Important pretraining parameters
A.6 Experimental Setup for Benchmarks
All fine-tuning experiments employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=1e-7$. Experiments were conducted on a single NVIDIA H100 GPU with a batch size of 500. When the full batch exceeded GPU VRAM, gradient accumulation was used. The learning rate schedule combined cosine annealing with a linear warmup: fine-tuning lasted 2000 steps if not specified differently, with the first 100 steps as warmup, and a minimum learning rate of 1e-8. During fine-tuning, each sample was augmented with two transformations—random rotations and random reflections.
A.6.1 MD17
For finetuning on the MD17 experiments we follow the experimental setup of [58], namely we sample 9500/500/10000 conformations for train/val/test splits for each molecule and evaluate the force prediction error on the test set. The hyperparameters for each molecule are shown in table 9.
Table 9: Fine-tuning Hyperparameter for MD17
Molecule Train Split Size Val Split Size Test Split Size Learning Rate
Aspirin 9500 500 10000 $5\times 10^{-5}$
Ethanol 9500 500 10000 $5\times 10^{-5}$
Naphthalene 9500 500 10000 $5\times 10^{-5}$
Salicylic Acid 9500 500 10000 $5\times 10^{-5}$
For full training runs on MD17 we use the same splits as above but train the model for 250k steps and with a batch size of 250, otherwise using identical settings to pretraining on QCML (see Appendix A.5)
For the stability evaluation we run an MD simulation with a timestep of 0.5fs for 600k steps. We use a Nosé-Hoover thermostat set to 500K. To improve predictions, we use frame-averaging with $n=4$ samples per step. For more details, see [58].
A.6.2 MD17@CCSD(T)
For the MD17@CCSD(T) experiments we adapt the experimental setup of Chmiela et al. [20] and extend the evaluation for MD-ET with a scenario with less training samples. Table 10 shows the hyperparameters for the MD17@CCSD(T) experiments.
Table 10: Fine-tuning Hyperparameter for MD17@CCSD(T)
Train Split Size Molecule Fine-tuning Steps Test Split Size Learning Rate
1000 Samples Ethanol 2000 500 $5\times 10^{-5}$
Aspirin 2000 500 $5\times 10^{-5}$
Benzene 2000 500 $5\times 10^{-5}$
Malonaldehyde 2000 500 $5\times 10^{-5}$
Toluene 2000 500 $5\times 10^{-5}$
100 Samples Ethanol 400 500 $5\times 10^{-5}$
Aspirin 400 500 $5\times 10^{-5}$
Benzene 400 500 $5\times 10^{-5}$
Malonaldehyde 400 500 $5\times 10^{-5}$
Toluene 400 500 $5\times 10^{-5}$
A.6.3 Ko2020
For the dataset introduced by [53] we follow the general experimental setup of Unke et. al. [13]. We additionally randomly split a validation set from the test split to optimize the learning rate. The hyperparameters used for fine-tuning for each system are shown in table 11.
Table 11: Fine-tuning Hyperparameter for Ko2020
Molecule Train Split Size Val Split Size Test Split Size Learning Rate
C10H2/C10H+${}_{3}^{+}$ 9035 50 930 $5\times 10^{-4}$
Na8/Cl+${}_{8}^{+}$ 4493 50 480 $1\times 10^{-4}$
Ag+/${}_{3}^{+/-}$ 9930 50 1030 $5\times 10^{-5}$
Au2–MgO 4468 50 480 $5\times 10^{-4}$
A.7 Experimental setup for evaluation of energy conservation
For the evaluation of energy conservation we run NVE simulations for 1 million using a step size of 0.5fs. To improve predictions, we use frame-averaging with $n=16$ samples per step.
Appendix B Visualization of molecules
Figure 4: Visualizations of the molecular structures used in this work. In the upper half are the cumulenes (left) and the alkanes (right) with increasing carbon chain length. On the bottom left are the four molecules used from MD17: naphthalene, aspirin, salicylic acid, and ethanol. On the bottom right are the four systems studied by Ko et al. [53]: a gold dimer on a magnesium oxide surface, a silver trimer, a salt system, and $\text{C}_{10}\text{H}_{2}$ (corresponding $\text{C}_{10}\text{H}^{+}$ omitted).
Algorithm 1 Comparison between standard attention and triangular attention
  function ATTENTION($X:n\times d$)
   $Q,K,V\leftarrow\text{linear}(X).\text{chunk}(3)$
   $\tilde{A}\leftarrow\text{einsum}(id,jd\rightarrow ij,Q,K)$
   $A\leftarrow\text{softmax}(\tilde{A}/\sqrt{d},-1)$
   $O\leftarrow\text{einsum}(ij,jd\rightarrow id,A,V)$
   return $\text{linear}(O)$
  end function
  function TRI_ATTENTION($X:n\times n\times d$)
   $Q,K,V^{1},V^{2}\leftarrow\text{linear}(X).\text{chunk}(4)$
   $V\leftarrow\text{einsum}(ild,ljd\rightarrow iljd,V^{1},V^{2})$
   $\tilde{A}\leftarrow\text{einsum}(ild,ljd\rightarrow ilj,Q,K)$
   $A\leftarrow\text{softmax}(\tilde{A}/\sqrt{d},-1)$
   $O\leftarrow\text{einsum}(ij,iljd\rightarrow ijd,A,V)$
   return $\text{linear}(O)$
  end function